OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (0, 2) and a directrix at y = -2

OpenStudy (anonymous):

@RadEn

OpenStudy (mathmale):

Would you suppose that this parabola is a vertical or a horizontal one? How would you know?

OpenStudy (anonymous):

I honestly have no idea..

OpenStudy (mathmale):

The focus is a point: (0,2), and the directrix is a straight line: y=-2. Are you able to graph these on coordinate axes, using the Draw feature (below)?

OpenStudy (anonymous):



OpenStudy (mathmale):

Very nice. The vertex of a parabola is always exactly halfway between the focus and the directrix. What point is the vertex of this parabola?

OpenStudy (anonymous):

(0, 0)

OpenStudy (mathmale):

So, your vertex is at (0,0) and your focus is at (0,2). Does this tell you whether the parabola is horizontal or vertical?

OpenStudy (anonymous):

It's vertical

OpenStudy (anonymous):

I think..

OpenStudy (anonymous):

Wait, is it horizontal?? If so, is the answer y^2=8x?

OpenStudy (mathmale):

See http://www.purplemath.com/modules/parabola.htm Your parabola is a vertical one. Yes, p=2, so 4py=8y=x^2, and so y = ??

OpenStudy (anonymous):

y=1/8 x^2?